shasha12:

Christian reads 1/4 book every 2/3 week. How many books does Christian read per week? ​ 1/6 ​ ​ 3/8 ​ ​ 2 2/3 ​ 6

shasha12:

3/8?

loveboy:

correct

shasha12:

thank you, how did you do that?

loveboy:

\[\large \bf \frac{2}{3} weeks= \frac{1}{4} book\] \[\large \bf 1~week=\frac{3 \times 1}{4 \times 2}book\] \[\large \bf 1~week=\frac{3}{8}book\]

shasha12:

thank you so much

loveboy:

your welcome